In a poker hand consisting of 5​ cards, find the probability of holding​ (a) 3 kings​; ​(b) 4 red cards and 1 club.


1
Expert's answer
2021-11-03T12:23:50-0400

for 52 card deck:


a)

"P=\frac{4}{52}\cdot \frac{3}{51}\cdot \frac{2}{50}\cdot \frac{48}{49}\cdot \frac{47}{48}=\frac{47}{270725}"


b)

"P=\frac{1}{2}\cdot \frac{25}{51}\cdot \frac{24}{50}\cdot \frac{23}{49}\cdot \frac{13}{48}=\frac{299}{19992}"